0 0
Add a comment Improve this question Transcribed image text
Answer #1

we know that range of ball thrown at an angle ? with initial velocity is given by
-0-sin 2θ R

therefore we know R=172m, ?=32.1 and g=9.8m/s2

substituting in the formula we get 172=v02 sin(64.2)/9.8

=>
v02  = 172x9.8/0.9 = 1872.22 => v0 =  43.27 m/s

therefore quarterback must throw the ball at a speed of 43.27 m/s
Answer #2

range = v0^2 sin(2 theta) /g

172 = v^2* sin (2*32.1)/9.9

172 = v^2 (sin64.2) /9.8

169.1 = v^2 * 0.90/9.8

v^2 =172 / 0.0918

v^2 = 1873.63
v = 43.28 m/s